OpenStudy (anonymous):

Please help me simplify 20z/sqrt(z^13).

OpenStudy (anonymous):

20z/sqrt((z^12)(z)) s0 20z/z^6(sqrt(z)) 20/z^5(sqrt(z)

OpenStudy (anonymous):

Thanks you, what does it mean in the second line where it says s0?

OpenStudy (anonymous):

sorry i want to write: so

OpenStudy (anonymous):

Oh, ok lol.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Thanks you!

OpenStudy (anonymous):

your welcome

OpenStudy (phi):

\[ \frac{20z}{\sqrt{z^{13}} } \\ \frac{20z}{\sqrt{z^{12}}\sqrt{z}}\\ \frac{20z}{z^{6}\sqrt{z}}\\ \] notice the z divided by sqr(z) is sqr(z) if you use exponents: \[ \frac{z}{z^\frac{1}{2}}= z^{1-\frac{1}{2}}= z^\frac{1}{2}=\sqrt{z} \] your problem becomes \[ \frac{20 \sqrt{z}}{z^6} \] which can be re-written a number of ways...

OpenStudy (anonymous):

Hmmm, ok.

OpenStudy (anonymous):

@phi I am a little confused. This question is from a multiple choice worksheet, and both your answer and sara17's answers are options. Which one should I use? Is one further simplified then the other?